Translate the following mathematical statements into mathematical equations and solve the unknown.
1. The current l varies directly as the electromotive force E and inversely as the resistance R. If a current l of 30 amperes flows through a system with 16 ohms resistance R and electromotive force E of 120 volts, find the current l when the system with 20 ohms resistance R and electromotive force E of 200- volt.
Solution 1
Step 1: Translate mathematical statement into mathematical equation
Step 2: Find the constant of variation.
Step 3: Give the equation of the variation.
Step 4: Find the value of unknown variable.

Answer by Boreal(15235)

I=k*E/R
30=k*120/16=7.5k
k=4
-
use that in the next part.
note that 2/3 again as much voltage and 20% more resistance, expect more current.
I=4*200/20
=40 amperes